OpenStudy (anonymous):

Caitlin had $402.00 in her bank account. She withdrew $15.00 each week to pay for a swimming lesson. She now has $237.00 in her bank account. a. Write an equation that can be used to find the number of swimming lessons that Caitlin paid for. b. How many swimming lessons did Caitlin pay for? c. At the time Caitlin had $237.00, the cost of a lesson rose to $19.00. How many lessons can Caitlin pay for with her remaining $237.00?

jimthompson5910 (jim_thompson5910):

let x = # of lessons she paid for

jimthompson5910 (jim_thompson5910):

if she started with $402 and now has $237, then she took out 402-237 = 165 dollars total So she spent $165 on x swimming lessons

jimthompson5910 (jim_thompson5910):

each lesson is $15, so this means 15x = 165

jimthompson5910 (jim_thompson5910):

to find part b), you just solve for x

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

or 237 + 15p = 402, where p = number of lessons paid for

OpenStudy (anonymous):

for all the problems this is the equation (starting amount of money) - (x)(price of lesson) = final amount of money x is the number of lessons.